A certain resistance thermometer reads 14.5 Ω in pure melting ice and 18.5 Ω in steam at standard atmospheric pressure. What will be its resistance at room temperature of 27.0oC?

1
Expert's answer
2022-03-24T16:18:54-0400
"R(t_1)=R_0(1+\alpha t_1)\\
R(t_2)=R_0(1+\alpha t_2)""R(27{^\circ C})=14.5\left(1+\frac{27}{100}*(18.5/14.5-1)\right)\\
=15.6\:\Omega"